OpenStudy (anonymous):

Step-by-Step How do i solve 4x^(2)-5x+5

OpenStudy (anonymous):

it has no real roots

OpenStudy (turingtest):

quadratic formula, there is no step by step to show, just plug in a=4 b=-5 c=5 if you don't know the quadratic formula look it up and memorize it, I don't feel like typing it

myininaya (myininaya):

is there something to solve here?

myininaya (myininaya):

i don't see anything to solve

OpenStudy (turingtest):

oh yeah, is it 4x^(2)-5x+5=0 ? otherwise myininaya is right should've noticed that

OpenStudy (turingtest):

@joyce you have to put an equals sign if you want us to do anything other than simplify, which in this case cannot be done be sure to type the entire problem, it makes all the difference

OpenStudy (anonymous):

this parabola opens up and vertex is above x-axis

OpenStudy (turingtest):

IF it is 4x^(2)-5x+5=y

OpenStudy (anonymous):

well the Q is asking me to factor 4x^(2)-5x+5.

OpenStudy (agreene):

\[x=\frac18(5\pm\sqrt{16y-55})\]

OpenStudy (turingtest):

can't be done

myininaya (myininaya):

not over the integers anyways

OpenStudy (agreene):

Oh, it can be done... just not over the reals.

OpenStudy (anonymous):

since the discriminant \[b ^{2}-4ac = 25 - 4(4)(5) = -55\]

OpenStudy (turingtest):

right, of course

myininaya (myininaya):

so it can be factored over the complex do you guys want to see? i can factor this by grouping

OpenStudy (turingtest):

yeah, let's see it

OpenStudy (agreene):

\[\frac{1}{16} (-8 i x+\sqrt{55}+5 i) (8 i x+\sqrt{55}-5 i)\]

OpenStudy (turingtest):

there it is, eh?

OpenStudy (agreene):

That's the lowest form I can find, at least.

OpenStudy (agreene):

But, I suck at factoring :D

OpenStudy (turingtest):

let's see what myininaya gets

myininaya (myininaya):

\[ax^2+bx+c=4x^2-5x+5=> a=4; b=-5;c=5\] \[b=(\frac{-5}{2}+z)+(\frac{-5}{2}-z)\] \[a \cdot c=(\frac{-5}{2}+z)(\frac{-5}{2}-z)\] We need to find what z remember that a*c=4(5)=20 so we can solve for z \[20=\frac{25}{4}-z^2 => 20-\frac{25}{4}=-z^2 => z^2=\frac{25}{4}-20\] \[z^2=\frac{25}{4}-\frac{20 \cdot 4}{4}=\frac{25-80}{4}=\frac{-55}{4}\] \[z=\pm \sqrt{\frac{-55}{4}}=\pm i \frac{\sqrt{55}}{2}\] We will only need \[z=i \frac{\sqrt{55}}{2}\] so we have \[bx=(\frac{-5}{2}+i \frac{\sqrt{55}}{2})x+(-\frac{5}{2}-i \frac{\sqrt{55}}{2})x\] so we have \[4x^2-5x+5\] now we will replace -5x as follows \[4x^2+(\frac{-5}{2}+i \frac{\sqrt{55}}{2})x+(\frac{-5}{2}-i \frac{\sqrt{55}}{2})x+5\] Now the factor by grouping happens! :) \[4x(x+\frac{-5+i \sqrt{55}}{8})+\frac{-5 - i \sqrt{55}}{2}(x+5 \cdot \frac{2}{-5 - i \sqrt{55}})\] Believe it or not we are almost there Now we will multiply top and bottom of 2/(-5-i sqrt{55}) by the conjugate of bottom \[4x(x+\frac{-5+i \sqrt{55}}{8})+\frac{-5-i \sqrt{55}}{2}(x+5 \cdot \frac{2(-5+i \sqrt{55})}{25-i^2 \cdot 55})\] \[4x(x+\frac{-5 + i \sqrt{55}}{8})+\frac{-5 - i \sqrt{55}}{2}(x+5 \cdot \frac{2(-5+i \sqrt{55})}{80})\] \[4x(x+\frac{-5 + i \sqrt{55}}{8})+\frac{-5- i \sqrt{55}}{2}(x+\frac{-5 + i \sqrt{55}}{8})\] Notice we have a common factor in each term \[(x+\frac{-5 + i \sqrt{55}}{8})(4x+\frac{-5 - i \sqrt{55}}{2})\]

OpenStudy (turingtest):

No, I just now finished reading both your link and proof and studied it closely, so it took a while to really get each step. I've never seen that before and really thought it was awesome, and saved it to my bookmarks. Thanks for showing me that it I really like to know how to do those tricky algebra things, they come in handy so often. Really epic, thanks! :D

OpenStudy (turingtest):

in the link? no way!

OpenStudy (turingtest):

Wow, very impressive. May I ask what you do or where you study?

OpenStudy (turingtest):

yeah, truly awesome :)

OpenStudy (turingtest):

Well, I knew you were pretty good but... wow, thanks.

OpenStudy (turingtest):

:)

OpenStudy (turingtest):

I have to assume you're a math major. Post Grad?

OpenStudy (alfie):

Great stuff, I enjoyed it.

OpenStudy (turingtest):

haha... yeah, I figured you were working with thesis-level material here.

OpenStudy (turingtest):

No, I just mean you're level was obvious. I'm sure you're not going to post your thesis on open study anytime soon.

OpenStudy (turingtest):

your*

OpenStudy (turingtest):

There's always a lot to learn.

OpenStudy (agreene):

Myiniaya I think I've seen something similar to your methodology you posted... it's kind of like the shortcut for complex factorization, only you used the quadratic as your mode... I'll look around my stuff and see if I can pull up the one that's similar... might take a bit, I think I learnt it in complex analysis as an undergrad >.<

OpenStudy (turingtest):

I'm just a first year physics major who has studied as much as I can online, so I can't relate to much high-level material. I do like differential equations though, and would consider looking more into PDF theory.

OpenStudy (turingtest):

PDE* partial differential equations typo

OpenStudy (agreene):

Most people have to take at least one physics course for their undergrad years

OpenStudy (turingtest):

It seems most have taken at least up to electromagnetism.

OpenStudy (turingtest):

you're physics teacher thought it was unfair for him to take math classes?

OpenStudy (turingtest):

I mean THE physics teacher, sorry

OpenStudy (turingtest):

but yeah I would hope so

OpenStudy (agreene):

The only paper I found from my notebook I thought it would be in was a proof that the roots of a polynomial can be multiplied to give the factorization. Which was how I did this problem... completed the square for roots and then multiplied them... more or less, the coeffs could be factored back in before the multiplication, so I did that, too. I'll keep looking around, I coulda sworn I saw something very similar >.<

OpenStudy (turingtest):

well that's partly why I tutor people here in math more than in physics.

OpenStudy (agreene):

lol I was exposed to vectors via math years before I started taking physics courses... made that bit much easier ;)

OpenStudy (turingtest):

just the opposite for me.

OpenStudy (turingtest):

I learned vectors in a physics class

OpenStudy (agreene):

I had to deal with vectors back in high school algebra O_o

OpenStudy (turingtest):

how's that possible

OpenStudy (turingtest):

yeah, that's all I learned in high school. Now I can do basic vector calculus and a fair amount of linear algebra .

OpenStudy (turingtest):

that's about all I know of vectors

OpenStudy (turingtest):

I had AP though

OpenStudy (agreene):

Yeah, we only did basic stuff... but it's things I see people learning in physics 1. Of course, by the time I was a senior we were doing basic vector calculus and in my non-euclidian geometry class we dealt with the problems that come up with vectors in those spaces.

OpenStudy (agreene):

...ever since that class i've been in love with projected geometry... Also, I live in SW Missouri, so you arkansas people were always being made fun of :P

OpenStudy (turingtest):

lol, so funny you seem to embrace your stereotype.

OpenStudy (turingtest):

Yeah, so where are they all smart? You're gonna say Boston, right?

OpenStudy (turingtest):

That's why I live in Mexico!

OpenStudy (agreene):

Projective Geometry is fun, I promise: Just look at how spheres are handled (about 1/2 way down the page) http://www.nct.anth.org.uk/basics.htm

OpenStudy (turingtest):

yup University of Gudalajara

OpenStudy (agreene):

I've been thinking about taking some classes at Universidad Autonoma de Yucatan (Merida)... fell in love with Merida when I first went there... god over a decade ago :X

OpenStudy (turingtest):

I'm trying to get to UNAM eventually, but I just got here and first need to get my papers in order... and a few credits under my belt too.

OpenStudy (agreene):

UNAM has an awesome library.

OpenStudy (agreene):

http://bnm.unam.mx/ and such.

OpenStudy (turingtest):

That I didn't know; yet another reason to go.

OpenStudy (turingtest):

thanks